OpenStudy (anonymous):

A group of private investors purchased a condominium complex for $4 million. They made an initial down payment of 12% and obtained financing for the balance. If the loan is to be amortized over 12 years at an interest rate of 13% per year compounded quarterly, find the required quarterly payment.

OpenStudy (anonymous):

$4million-12%*4million(1+.13)^48/[(1+.13)^48-1]

OpenStudy (anonymous):

$ to be financed = 0.88*4000000 or 4 millon = $3,520,000 r = 0.0325/qr (13%/4) , n = 48 qrs PMT = PVr/(1-(1+r)^-n) = 3,520,000*.0325/(1-(1.0325^-48)) = $145,809.6783